# Forecasting Automotive Supply Chain Disruption with Heterogeneous Time Series

### v1

@@
Forecasting Automotive Supply Chain Disruption with Heterogeneous Time Series
@@
[…]
Operational disruptions can significantly impact a company’s performance. Ford, with its 37 plants globally, utilizes 17 billion parts annually to manufacture six million cars and trucks. Given the presence of up to ten tiers of suppliers between the company and the raw materials, any extended disruption within this extensive supply chain can inflict substantial financial losses. Therefore, the ability to forecast and identify such disruptions early is crucial for maintaining seamless operations. In the contemporary era of Big Data, information on suppliers’ past performance has become an invaluable resource for analyzing and predicting disruption risks. In this study, we demonstrate how we construct a dataset consisting of many multivariate time series to forecast first-tier supply chain disruptions, utilizing features related to capacity, inventory, utilization, and processing, as outlined in the classical Factory Physics framework. From a technical standpoint, this dataset presents significant challenges due to its complexity and vast scale, encompassing over five hundred thousand individual series. Furthermore, these time series, while exhibiting certain similarities, also display heterogeneity within specific subgroups. As a result, traditional industrial and statistical models are inadequate for capturing the complex dynamics and managing the extensive volume of data. To address these challenges, we propose a novel methodology that integrates an enhanced Attention Sequence to Sequence Deep Learning architecture, using Neural Network Embeddings to model group effects, with a Survival Analysis model. This model is designed to learn intricate heterogeneous data patterns related to operational disruptions. Our model has demonstrated a strong performance, achieving 85% precision and 80% recall during the Quality Assurance (QA) phase across Ford’s five North American plants. While originally developed for supply chain time series data, our methodology can be generalized to any dataset comprising multivariate time series with inherent heterogeneity. Additionally, to address the common criticism of Machine Learning models as ”black boxes,” we show how the SHAP (SHapley Additive exPlanations) framework can be used to generate feature importance from the proposed model’s predictions. It offers valuable insights that can lead to actionable strategies and highlights the potential of advanced machine learning techniques for managing and mitigating supply chain risks in the automotive industry.
[…]
Supply Chain Disruption; Supply Chain Resilience; Supply Chain Deep Learning; Survival Analysis; Sequence-to-Sequence
@@ 1 Introduction
Ford maintains a complex supply chain and operational network, operating 37 plants globally, utilizing 17 billion parts annually to manufacture six million cars and trucks. The company has up to 10 tiers of suppliers between itself and raw materials. An extended disruption anywhere within this extensive supply chain can inflict a substantial financial impact on the company. In the literature, scholars and practitioners generally agree that disruption negatively impacts the company (sheffi2005; hendricks2005; network2013building). However, there is less agreement on classifying and forecasting such disruption (kleindorfer2005; tang2006; wagner2006; sodhi2012). Understanding the different sources and risks of disruption is critical because we can make informed decisions on which disruptions warrant mitigation investment (simchi2014). Supply chain disruptions can stem from many sources, such as global pandemics, natural disasters, geopolitical risks, terrorist attacks, environmental hazards, volatile fuel prices, rising labor costs, currency fluctuations, counterfeit parts and products, delivery delays, market changes, and supplier performance issues.
@@
In the supply chain literature, efforts have been dedicated to modeling the propagation of disruption effects through supply chain networks. One of the early studies by hopp2006 analyzed the propagation of disruption effects through a simplified supply chain. schmitt2011 evaluated response strategies to minimize service level impacts in a multi-echelon network during random-duration disruptions. mackenzie2014 examined the interaction between supplier and buyer response strategies under random-duration disruptions. Simchi-Levi et al. (simchi2014; simchi2015) introduced the Time-To-Recover (TTR) metric to quantify the financial impact of disruptions on the entire supply chain, measured by the Risk Exposure Index (REI). This index enables companies to rank their direct and indirect suppliers, identifying the ”weak links” in their supply chains. Additionally, simchi2015 proposed the Time-To-Survive (TTS) concept, which defines the maximum duration the entire supply chain can generally function before the ripple effects of a disruption impact performance. These concepts—TTR, REI, and TTS—have been implemented at Ford Motor Company to manage supply chain risks (simchi2015).
@@
These studies are based on the assumption that disruption events can be identified well in advance. In the era of Big Data, the rise of Machine Learning and Artificial Intelligence offers opportunities for predicting disruptive events with high degrees of accuracy. However, the literature on utilizing Machine Learning for supply chain risk management remains sparse. fan2015 explored the potential of leveraging big data sources related to supply chains and proposed a Supply Chain Risk Management (SCRM) framework to detect emergent risks. he2020 recognized the predictive capabilities of incorporating a significant data analytical component into a generic SCRM framework. Nevertheless, these works are theoretical and lack real-world application or implementation of the proposed frameworks and models. More concretely, ye2015 used a Support Vector Machine classifier to identify disruptions based on the economic performance of firms within the supply chain, collecting public financial data for these firms before, during, and after supply chain disruptions. brintrup2020 analyzed historical data from an Original Equipment Manufacturer (OEM) that produced complex engineering assets to predict order delays using machine learning models, including Random Forest, Support Vector Machine, Logistic Regression, and Linear Regression. In their study, the Random Forest Classifier achieved the best performance. Although both studies identify potential suppliers with a high risk of disruptions, they do not provide estimates for the time until these disruption events occur. Such estimates are crucial for integrating this research into the broader framework of Supply Chain Risk Management.
As a leading automobile manufacturer, Ford can request and accumulate an extensive repository of proprietary data on suppliers’ capacity and performance, providing a potent source for highly accurate predictive capabilities. Moreover, Deep Learning has recently emerged as preeminent models in Artificial Intelligence, driven by a decade of rigorous research. Consequently, many state-of-the-art Machine Learning models are now based on Deep Learning.
@@
Our contribution in this work is to advance the research on applying Big Data to forecast supply chain disruptions. We demonstrate the construction of a complex dataset comprising many multivariate time series that track the arrangements between Ford and our first-tier suppliers for transporting critical vehicle parts to Ford’s manufacturing plants. We meticulously select data features that reflect capacity, inventory, utilization, and process time—key aspects in the classical Factory Physics (see hopp2011factory). Furthermore, we propose an AI model that integrates an enhanced Sequence to Sequence with Attention Deep Learning architecture with a parametric survival analysis likelihood model. This enhanced architecture is designed to model the heterogeneity in the data, driven by various combinations of plants, sites, and vehicle parts. Our work introduces an important improvement to the Seq2Surv model proposed by li2022, developed for generic time series exhibiting similar underlying statistical patterns despite random variations. In contrast, the disruption behaviors in our data can differ markedly, even contradictorily, based on the specific values of suppliers, plants, and parts. While originally developed for supply chain time series data, our methodology can be generalized to any dataset comprising multivariate time series with inherent heterogeneity. Lastly, we illustrate how to use the SHAP (SHapley Additive exPlanations, phillips2021) framework to calculate feature importance for each of the model’s predictions. This explainable AI technique provides valuable insights that can inform actionable strategies for our business partners.
@@
Our model relies on the Survival Analysis framework to model observed time-to-disruption data. Survival Analysis, a branch of Statistics, focuses on the study of time-to-event data. This field encompasses a variety of applications, such as predicting the survival of cancer patients (vigano2000), customer churn (van2004), mechanical system failure (susto2014), credit scoring (dirick2017), and reliability and manufacturing problems (li2022). The strength of Survival Analysis lies in its interpretability, flexibility, and ability to handle censored data. However, one of its notable weaknesses is predictive accuracy. To enhance predictive performance, numerous studies have extended classical Survival Analysis using Neural Networks.
@@
For discrete-time survival problems, lee2018 applied neural networks to the discrete-time likelihood for right-censored time-to-event data, parameterizing the probability mass function. gensheimer2019 adopted a similar likelihood approach, parameterizing hazard rates with a neural network. kvamme2019discrete, through simulation studies and real-world data, found that hazard rate parameterization performed slightly better. Building on this insight, the authors introduced PC-Hazard, which parameterization the hazard rate for continuous survival time data by discretizing the continuous time scale, assuming the continuous-time hazard is piece-wise constant.
The remainder of the paper is organized as follows: Section 2 reviews the relevant background and preliminaries. Section 3 offers a detailed explanation of how to construct a dataset for forecasting supply chain disruptions by selecting features that reflect key aspects of classical Factory Physics. In Section 4, we present the Heterogeneous Sequence-to-Disruption AI model. Section 5 evaluates the model’s performance and discusses methods for interpreting the results using the SHAP framework. Finally, Section 6 concludes the paper.
@@ 2.1 Survival Analysis
In this study, we aim to model the distribution of the time-to-disruption, denoted as $T^{*}$. In practical scenarios, not all disruption times are observable and are often subject to right censoring, wherein the observation period concludes before the disruption event occurs. Let $C^{*}$ represent the time to the censoring event. Formally, we define:
@@
Here, $T$ is the observed time-to-disruption, and $Y$ is an indicator that equals $1$ if the disruption event occurs within the observation window $T^{*}\leq C^{*}$ and 00 otherwise. This framework is fundamental in Survival Analysis, a branch of statistics focusing on time-to-event data (jenkins2005). As demonstrated in the next section, in our data, the time-to-disruption is discrete, ranging from $1$ to $365$ days. The discrete disruption time can be modeled using a parametric approach (see Chapter 10 in moore2016).
Before we define the model, let’s go over some preliminaries. Consider a random discrete variable $T^{*}$ with the discrete support $\tau_{1}<\tau_{2}<\ldots<\tau_{j}<\ldots$, the cumulative distribution function (CDF) is given by $F(t)=P(T^{*}\leq t)=\sum_{i=0}^{t}\mathbb{P}(T^{*}=t)$, and the corresponding probability mass function (PMF) is $f(t)=\mathbb{P}(T^{*}=t)$.
In Survival Analysis, it’s often more useful to discuss the survival function, defined as $S(t)=1-F(t)=P(T^{*}>t)$, along with its associated hazard function. For a discrete variable $T^{*}$ and two consecutive survival time points $\tau_{j-1}<\tau_{j}$, the hazard function is denoted $h(\tau_{j})=\mathbb{P}(T^{*}=\tau_{j}\;|\;T^{*}>\tau_{j-1})=\cfrac{f(\tau_{j})}{S(\tau_{j-1})}=\cfrac{S(\tau_{j-1})-S(\tau_{j})}{S(\tau_{j-1})}$. As such,
@@
$\mathbb{P}(T=t,Y=y)$ $=\mathbb{P}(T=t,t\leq C^{*})^{y}\mathbb{P}(T=t,t>C^{*})^{1-y}$
@@
The third equality is due to $T^{*}$ and $C^{*}$ are independent. Since random variable $T$ is discrete, we can model the hazard function using a parametric distribution with parameter $\theta$. We define $\kappa(t)$ as the index of the discrete time point in the support, i.e., $t=\tau_{\kappa(t)}$. For a sample of observations $(t_{1},y_{1}),(t_{2},y_{2}),\ldots,(t_{n},y_{n})$, the log-likelihood function can be expressed as a sum of individual log-likelihood contributions as follows using equations (1) and (2).
@@
In the second equality, terms related to censoring times $C_{i}$ are omitted from the log-likelihood function as they are constant and do not contain parameter $\theta$. Utilizing equation 2 recursively in the final term of the last equality above, we can finally write the log-likelihood of the sample data as,
@@
Observe that the equation 3 consists solely of the hazard function. To model the observed data, we simply need to define the form of the hazard function. Since the hazard function represents a probability in discrete form, its range must lie between 0 and 1. One suitable choice is the logistic hazard function (see brown1975). Assume that the time-to-disruption is also conditioned on covariates/features $X$,
@@ 2.2 Sequence-to-Sequence Architecture
A significant enhancement to the Seq2Seq model is integrating the Attention mechanism proposed by bahdanau2014. This mechanism enables the model to dynamically focus on different parts of the input sequence while generating each output unit. Specifically, the context vector is computed as a weighted sum of the encoder’s hidden states, with the weights reflecting the relevance of each hidden state to the current decoding step. This dynamic focusing capability significantly improves the model’s performance, particularly for tasks involving long and complex input sequences.
@@ 2.3 Seq2Surv Model
Leveraging the power of Sequence-to-Sequence neural networks and the Attention mechanism for modeling sequential data, li2022 employs this advanced deep learning architecture to handle time series data in the reliability and manufacturing domain. This approach replaces the linear predictor in Cox proportional hazard models (see Chapter 5 in moore2016) and discrete time logistic hazard parametric models (see equation 4) with learnable non-linear functions. In this methodology, time series are treated as sequential input to the Encoder, which li2022 implemented using Bidirectional Gated Recurrent Units (GRUs), as introduced by cho2014. The GRU, similar to a Long Short-Term Memory network (LSTM, see hochreiter1997) with its gating mechanisms for input and forgetting features, lacks output gates, resulting in fewer parameters than LSTMs. The Decoder in Seq2Surv then generates a sequence of survival probability estimates for the entire lifetime, which is then used in the Survival Analysis’ log-likelihood function to model the observed disruption times.

### v3

# Forecasting Automotive Supply Chain Shortfalls with Heterogeneous Time Series

@@
Forecasting Automotive Supply Chain Shortfalls with Heterogeneous Time Series
@@
Operational shortfalls can have a major impact on a company’s performance. Ford, with its extensive network of plants around the world, uses tens of billions of parts each year to manufacture millions of cars and trucks. With up to ten levels of suppliers between the company and raw materials, any extended disruption in this supply chain can lead to substantial financial losses. Therefore, the ability to predict and identify early shortfalls is crucial to maintaining seamless operations. In this study, we demonstrate how we constructed a dataset of multivariate time series to forecast first-tier supply chain shortfalls, using features related to capacity, variability, utilization, and processing time. This dataset is technically challenging due to its vast scale of over five hundred thousand time series. Furthermore, these time series, while exhibiting certain similarities, also display heterogeneity within specific subgroups. To address these challenges, we propose a novel methodology that integrates an enhanced Attention Sequence to Sequence Deep Learning architecture, using Neural Network Embeddings to model group effects, with a Survival Analysis model. This model is designed to learn intricate heterogeneous data patterns related to operational shortfalls. Our model has demonstrated a strong performance, achieving 0.85 precision and 0.8 recall during the Quality Assurance (QA) phase across Ford’s five North American plants. Furthermore, to address the common criticism of deep learning models as “black boxes,” we show how the SHAP framework can be used to generate the importance of characteristics from the predictions of the proposed model. It offers valuable insights that can lead to actionable strategies and highlights the potential of advanced machine learning for managing and mitigating supply chain risks in the automotive industry. 111 The information presented in this paper is intended solely for academic research purposes and should not be construed or used as financial advice.
@@
Supply Chain Shortfalls; Supply Chain Resilience; Supply Chain Deep Learning; Survival Analysis; Sequence-to-Sequence
@@ 1 Introduction
Ford maintains a complex supply chain and operational network, operating 37 plants globally, utilizing more than 17 billion parts annually to manufacture about six million cars and trucks. The company has up to 10 tiers of suppliers between itself and raw materials. An extended disruption anywhere within this extensive supply chain can inflict a substantial financial damage on the company. In the literature, scholars and practitioners generally agree that disruption negatively impacts the company (sheffi2005; hendricks2005; network2013building). However, there is less agreement on classifying and forecasting such disruptions (kleindorfer2005; tang2006; wagner2006; sodhi2012). Understanding the different sources and risks of disruption is critical because we can make informed decisions on which disruptions warrant mitigation investment (simchi2014).
A supply chain disruption refers to any incident that adversely affects the production, distribution, or sales of products. Such disruptions can stem from a wide array of sources, including global pandemics, natural disasters, geopolitical risks, terrorist attacks, environmental hazards, fluctuating fuel prices, increasing labor costs, currency fluctuations, counterfeit parts and products, delivery delays, market changes, and issues related to supplier performance. Our study focuses on forecasting the potential disruptions where the associated risks are known and controllable. Specifically, our work at Ford aims to forecast the time until occurrences where assembly parts delivery is being delayed beyond their scheduled production date at Ford assembly plants. Throughout this paper, we refer to these events as “shortfall events.”
@@
In the supply chain literature, efforts have been dedicated to modeling the propagation of disruption effects through supply chain networks. One of the early studies by hopp2006 analyzed the propagation of disruption effects through a simplified supply chain. schmitt2011 evaluated response strategies to minimize service level impacts in a multi-echelon network during random-duration disruptions. mackenzie2014 examined the interaction between supplier and buyer response strategies under random-duration disruptions. Simchi-Levi et al. (simchi2014; simchi2015) introduced the Risk Exposure Index (REI) model to quantify the impact of disruptions at a given supplier. This index enables companies to rank their direct and indirect suppliers, identifying the ”weak links” in their supply chains. Additionally, simchi2015 proposed the Time-To-Survive (TTS) concept, which defines the maximum duration the entire supply chain can generally function before the ripple effects of a disruption impact performance. Ripple effect (see ivanov2019handbook) describes the impact of a disruption propagation on supply chain performance as well as scope of changes in the structural design and planning parameters of the supply chain. The concepts—TTR, REI, and TTS—have been implemented at Ford Motor Company to manage supply chain risks (simchi2015). More recently, dolgui2024network proposed and analyzed a ”network-of-networks” mechanism as a cross-industry adaptation strategy to mitigate uncertainties in supply chains during crises such as the COVID-19 pandemic.
@@
In the era of Big Data, the rise of Machine Learning and Artificial Intelligence offers new opportunities to leverage massive amount of data to predict these disruptive events. Nevertheless, the literature on utilizing Machine Learning for supply chain risk management remains sparse. fan2015 explored the potential of leveraging big data sources related to supply chains and proposed a Supply Chain Risk Management (SCRM) framework to detect emerging risks. he2020 recognized the predictive capabilities of incorporating a significant data analytical component into a generic SCRM framework. Nevertheless, these works are theoretical and lack real-world application or implementation of the proposed frameworks and models. More concretely, ye2015 used a Support Vector Machine classifier to identify disruptions based on the economic performance of firms within the supply chain, collecting public financial data for these firms before, during, and after supply chain disruptions. brintrup2020 analyzed historical data from an Original Equipment Manufacturer (OEM) that produced complex engineering assets to predict order delays using machine learning models, including Random Forest, Support Vector Machine, Logistic Regression, and Linear Regression. In their study, the Random Forest Classifier achieved the best performance. Although both studies identify potential suppliers with a high risk of disruptions, they do not provide estimates for the time until these disruption events occur. Such estimates are crucial for integrating this research into the broader framework of Supply Chain Risk Management.
@@
As a leading automobile manufacturer, Ford can collect an extensive repository of historical data on suppliers’ capacity and performance, providing a potent source for highly accurate predictive capabilities. Deep Learning has recently emerged as preeminent models in Artificial Intelligence, driven by a decade of rigorous research. Consequently, many state-of-the-art Machine Learning models are now based on Deep Learning. Our contribution in this work is to advance the research on applying Big Data to forecast supply chain shortfall events with sophisticated Deep Learning models. We demonstrate the construction of a complex dataset comprising many multivariate time series that track the arrangements between Ford and our first-tier suppliers for transporting critical vehicle parts to Ford’s manufacturing plants. We meticulously select data features that reflect capacity, inventory, utilization, and process time, key aspects in the classical Factory Physics (see hopp2011factory). Furthermore, we propose an AI model that integrates an enhanced Sequence to Sequence with Attention Deep Learning architecture with a parametric survival analysis likelihood model. This enhanced architecture is designed to model the heterogeneity in the data, driven by various combinations of plants, sites, and vehicle parts. Our work introduces an important improvement to the Seq2Surv model proposed by li2022, developed for generic time series that exhibit similar underlying statistical patterns despite random variations. In contrast, the shortfall behaviors in our data can differ significantly, even contradictorily, based on the specific combinations of suppliers, plants, and parts. While originally developed for supply chain time series data, our methodology can be generalized to any dataset comprising multivariate time series with inherent heterogeneity. Lastly, we illustrate how to use the SHAP (SHapley Additive exPlanations, phillips2021) framework to calculate feature importance for each of the model’s predictions. This explainable AI technique provides valuable insights that can inform actionable strategies for our business partners.
@@
Our model relies on the Survival Analysis framework to model observed time-to-disruption data. Survival Analysis, a branch of Statistics, focuses on the study of time-to-event data (jenkins2005). This field encompasses a variety of applications, such as predicting the survival of cancer patients (vigano2000), customer churn (van2004), mechanical system failure (susto2014), credit scoring (dirick2017), and reliability and manufacturing problems (krivtsov2002; li2022). The strength of Survival Analysis lies in its interpretability, flexibility, and ability to handle censored data. However, one of its notable weaknesses is predictive accuracy. To enhance predictive performance, numerous studies have extended classical Survival Analysis using Neural Networks.
@@
For discrete-time survival problems, lee2018 applied neural networks to the discrete-time likelihood for right-censored time-to-event data, parameterizing the probability mass function. gensheimer2019 adopted a similar likelihood approach, parameterizing hazard rates with a neural network. kvamme2019discrete, through simulation studies and real-world data, found that hazard rate parameterization performed slightly better. Building on this insight, the authors introduced PC-Hazard, in which the parameterization of the hazard rate for continuous survival time data is achieved by discretizing the continuous time scale, assuming the continuous-time hazard is piecewise constant.
The remainder of the paper is organized as follows: Section 2 reviews the relevant background and preliminaries. Section 3 offers a detailed explanation of how to construct a dataset for forecasting supply chain disruptions by selecting informative features. In Section 4, we present the Heterogeneous Sequence-to-Disruption AI model. Section 5 evaluates the model’s performance and discusses methods for interpreting the results using the SHAP framework. Finally, Section 6 concludes the paper.
@@ 2.1 Survival Analysis
In this study, we aim to model the distribution of the time-to-shortfall, denoted as $T^{*}$. In practical scenarios, not all shortfall times are observable and are often subject to right censoring, wherein the observation period concludes before the shortfall event occurs. Let $C^{*}$ represents the time to the censoring event. Formally, we define:
@@
Here, $T$ is the observed time-to-shortfall, and $Y$ is an indicator that equals $1$ if the shortfall event occurs within the observation window $T^{*}\leq C^{*}$ and 00 otherwise. This framework is fundamental in Survival Analysis. As demonstrated in the next section, in our data, the time-to-shortfall is discrete, ranging from $1$ to $365$ days. The discrete disruption time can be modeled using a Survival Analysis parametric approach (see Chapter 10 in moore2016).
Before we define the model, we review some preliminaries. Consider a random discrete variable $T^{*}$ with the discrete support $\tau_{1}<\tau_{2}<\ldots<\tau_{j}<\ldots$, the cumulative distribution function (CDF) is given by $F(t)=P(T^{*}\leq t)=\sum_{i=0}^{t}\mathbb{P}(T^{*}=i)$, and the corresponding probability mass function (PMF) is $f(t)=\mathbb{P}(T^{*}=t)$ where $t$ is a positive integer.
In Survival Analysis, it’s often more useful to work with the survival function, defined as $S(t)=1-F(t)=P(T^{*}>t)$, along with its associated hazard function. For a discrete variable $T^{*}$ and two consecutive survival time points $\tau_{j-1}<\tau_{j}$, the hazard function is denoted $h(\tau_{j})=\mathbb{P}(T^{*}=\tau_{j}\;|\;T^{*}>\tau_{j-1})=\cfrac{f(\tau_{j})}{S(\tau_{j-1})}=\cfrac{S(\tau_{j-1})-S(\tau_{j})}{S(\tau_{j-1})}$. As such,
@@
$\mathbb{P}(T=t,Y=y)$ $=\mathbb{P}(T^{*}=t,t\leq C^{*})^{y}\mathbb{P}(T^{*}>t,C^{*}=t)^{1-y}$
@@
The first equality follows the definition of $T^{*}$ and $C^{*}$ under the events $\{Y=1\}$ and $\{Y=0\}$. The second equality is due to the independence of $T^{*}$ and $C^{*}$. The third equality also holds due to their independence. Because $T$ is discrete, we can model its hazard function using a parametric distribution with parameter $\theta$. We define $\kappa(t)$ as the index of the discrete time point in the support, i.e., $t=\tau_{\kappa(t)}$. For a sample of observations $(t_{1},y_{1}),(t_{2},y_{2}),\ldots,(t_{n},y_{n})$, the log-likelihood function can be expressed as a sum of individual log-likelihood contributions as follows using equations (1) and (2).
@@
In the second line, terms related to censoring times $C_{i}$ are omitted from the log-likelihood function as they are constant with respect to the parameter $\theta$. Utilizing equation 2 recursively in the final term of the last equality above, we can finally write the log-likelihood of the sample data as,
@@
Observe that the equation 3 consists solely of the hazard function. To model the observed data, we simply need to define the form of the hazard function. Since this hazard function represents a probability of discrete variable $T$, its range must lie between 0 and 1. One suitable modeling choice is the logistic hazard function (see brown1975). Assume that the time-to-shortfall is also conditioned on covariates/features $X$,
@@ 2.2 Sequence-to-Sequence Architecture
A significant enhancement to the Seq2Seq model is the integration of Attention mechanism proposed by bahdanau2014. This mechanism enables the model to dynamically focus on different parts of the input sequence while generating each output unit. Specifically, the context vector is computed as a weighted sum of the encoder’s hidden states, with the weights reflecting the relevance of each hidden state to the current decoding step. This dynamic focusing capability significantly improves the model’s performance, particularly for tasks involving long and complex input sequences.
@@ 2.3 Seq2Surv Model
Leveraging the power of Sequence-to-Sequence neural networks and the Attention mechanism for modeling sequential data, li2022 employed this advanced deep learning architecture to handle time series data in the reliability and manufacturing domain. This approach replaces the linear predictor in Cox proportional hazard models (see Chapter 5 in moore2016) and discrete time logistic hazard parametric models (see equation 4) with learnable non-linear functions. In this methodology, time series are treated as sequential input to the Encoder, which li2022 implemented using Bidirectional Gated Recurrent Units (GRUs), as introduced by cho2014. The GRU, similar to a Long Short-Term Memory network (LSTM, see hochreiter1997) with its gating mechanisms for input and forgetting features, lacks output gates, resulting in fewer parameters than LSTMs. The Decoder in Seq2Surv then generates a sequence of survival probability estimates for the entire lifetime, which is then used in the Survival Analysis’ log-likelihood function to model the observed disruption times.